Find the elasticity of sustitution for the production function Q=√LK

1
Expert's answer
2022-02-08T12:09:42-0500

Q = L0.5K

Elasticity of substitution ("\alpha") = "\frac{percentage change in K/L}{percentage change in MRTS}"


% change in K/L: = "{dIn(K\/L)}"

MRTS: ="\frac{MPL}{MPK}"


"MPL=\frac {\delta Q}{\delta L}=0.5L^{-0.5}K"


 "MPK=\frac{\delta Q}{\delta K }=L^{0.5}"


"MRTS=\frac{0.5L^{-0.5}K}{L^{0.5}}=\frac{0.5K}{L}"


"\alpha" = "\frac{dIn(0.5K/L)}{dIn(0.5K/L)}"


"\alpha" = 1